# The Unbounded Denominators Conjecture

### v1

@@ 1. Introduction
(1.0.2) $\lambda(\tau)=16q\prod_{n=1}^{\infty}\left(\frac{1+q^{2n}}{1+q^{2n-1}}\right)^{8}=16q-128q^{2}+\cdots$
with $q=e^{\pi i\tau}$. (Historic conventions force one to use $q$ for both $e^{\pi i\tau}$ and $e^{2\pi i\tau}$ — we use the first choice unless we expressly state otherwise.) On replacing the weight $k$ form $f$ by the weight zero form $f(\tau)^{12}(\lambda(\tau)/16\Delta(\tau/2))^{k}$, where $\Delta(\tau/2):=q\prod_{n=1}^{\infty}(1-q^{n})^{24}$, we may (and do) assume that $k=0$. The function $f$ is then an algebraic function of $\lambda$, with branching only at the three punctures $\lambda=0,1,\infty$ of the modular curve $Y(2)\cong\mathbf{P}^{1}\smallsetminus\{0,1,\infty\}$. Thus another reading of our result states that the Belyĭ maps (étale coverings)
@@
The resulting Fourier coefficients are closely linked to the dimensions of the irreducible representations of the exceptional Lie group $E_{8}(\mathbf{C})$, and in particular they are integers. To be more precise: the modular function $j^{1/3}$ coincides with the graded dimension of the level one highest-weight representation of the affine Kac–Moody algebra $E_{8}^{(1)}$; see Gannon’s book [Gan06, page 196] for a broad view on this topic and its relation to mathematical physics. The unbounded denominators conjecture (Theorem 1.0.1) now implies that $j^{1/3}$ must be a modular function on a congruence subgroup, and indeed one may easily confirm that it is a Hauptmodul for $\Gamma(3)$. Here the complete list of $n$ for which $j^{1/n}$ is congruence modular are the divisors of $24$. In a similar vein pertaining to the examples from the representation theory of vertex operator algebras, we prove in our closing section 7 the natural generalization of Theorem 1.0.1 to components of vector-valued modular forms for $\mathrm{SL}_{2}(\mathbf{Z})$, in particular resolving — in a sharper form, in fact — Mason’s unbounded denominators conjecture [Mas12, KM08] on generalized modular forms.
@@ 1.1.1. The Diophantine principle
We shall refer to the first inequality as a Liouville lower bound, following its use by Liouville in his proof of the lower bound $|\alpha-p/q|\gg 1/q^{n}$ for algebraic numbers $\alpha\neq p/q$ of degree $n\geq 1$. We shall refer to the second inequality as a Cauchy upper bound, following the example above where it comes from an application of the Cauchy integral formula. The first non-trivial generalization of Lemma 1.1.1 was Émile Borel’s theorem [Bor94]. Dwork famously used a $p$-adic generalization of Borel’s theorem in his $p$-adic analytic proof of the rationality of the zeta function of an algebraic variety over a finite field (see Dwork’s account in the book [DGS94, Chapter 2]). The simplest non-trivial statement of Borel’s theorem is that an integral formal power series $f(x)\in\mathbf{Z}\llbracket x\rrbracket$ must already be a rational function as soon as it has a meromorphic representation as a quotient of two convergent complex-coefficients power series on some disc $D(0,R)$ of a radius $R>1$. The subject of arithmetic algebraization blossomed at the hands of many authors, including most prominently Carlson, Pólya, Robinson, D. &\& G. Chudnovsky, Bertrandias, Zaharjuta, André, Bost, Chambert-Loir [CL02, BCL09, Ami75], [And04, § I.5], [And89, § VIII]. A simple milestone that we further develop in our § 2 is André’s algebraicity criterion [And04, Théorème 5.4.3], stating in a particular case that an integral formal power series $f(x)\in\mathbf{Z}\llbracket x\rrbracket$ is algebraic as soon as the two formal functions $x$ and $f$ admit a simultaneous analytic uniformization — that means an analytic map $\varphi:(D(0,1),0)\to(\mathbf{C},0)$ such that the composed formal function germ $f(\varphi(z))\in\mathbf{C}\llbracket z\rrbracket$ is also analytic on the full disc $D(0,1)$, and such that $\varphi$ is sufficiently large in terms of conformal size, namely: $|\varphi^{\prime}(0)|>1$. For example, for any integer $m$, the algebraic power series $f=(1-m^{2}x)^{1/m}\in\mathbf{Z}\llbracket x\rrbracket$ admits the simultaneous analytic uniformization $x=(1-e^{Mz})m^{-2}$ and $f=e^{Mz/m}$, where the conformal size $|\varphi^{\prime}(0)|=M/m^{2}$ can clearly be made arbitrarily large by making a suitable choice of $M$.
@@
In our setting, the Liouville lower bound ultimately comes down to the integrality (“bounded denominators”) hypothesis on the Fourier coefficients of $f(\tau)$, while the Cauchy upper bound comes down to studying the mean growth behavior $m(r,\varphi):=\int_{|z|=r}\log^{+}{|\varphi|}\,\mu_{\mathrm{Haar}}$ of the largest (universal covering) analytic map $\varphi:D(0,1)\to\mathbf{C}\smallsetminus\mu_{N}$ avoiding the $N$-th roots of unity. These are clearly distinguished in our abstract arithmetic algebraization work of § 2 as the steps (2.4.1) and (2.3.10), respectively. Our Theorem 2.0.2 is effectively a quantitative refinement of André’s algebraicity criterion to take into account the degree of algebraicity over $\mathbf{Q}(x)$, and still more precisely a certain holonomy rank over $\mathbf{Q}(x)$. Foreshadowing a key technical point (to be discussed in more detail later in the introduction), our Cauchy upper bound is given in terms of a mean (integrated) growth term rather than a supremum term, and this improvement is essential to our approach.
@@ 1.1.2. Modularity and simultaneous uniformizations of $f$ and $\lambda$
of (1.0.2) has integer coefficients, expressing the identity $\mathbf{Z}\llbracket q\rrbracket=\mathbf{Z}\llbracket x\rrbracket$ of formal power series rings, and that formal substitution turns our integral Fourier coefficients hypothesis into an algebraic power series with integer coefficients: henceforth in this introductory sketch we switch to writing, by a mild and harmless notational abuse, simply $f(x)\in\mathbf{Z}\llbracket x\rrbracket$ in place of $f(\tau)$ and $\lambda(q)$ in place of (1.0.2). In the general case of arbitrary cusp width, which we need anyhow for the inner workings of our proof even if one is ultimately interested in the $\mathbf{Z}\llbracket q\rrbracket$ case, we will only have $f\in\mathbf{Z}[1/N]\llbracket x\rrbracket$ — but there is still a hidden integrality property which we can exploit. That leads to some mild technical nuance with the power series (2.0.3) — think of $t=q^{1/N}$, $x(t)=\sqrt[N]{\lambda(t^{N})/16}$ and $p(x)=x^{N}$ — in our refinement (2.0.4) of André’s theorem.
The complex analysis enters by way of a linear ODE in the following way. To start with, we have, just by fiat, the simultaneous analytic uniformization of the two functions $x=\lambda/16$ and $f$ by the complex unit $q$-disc $|q|<1$. In this way, the tautological choice $\varphi(z):=\lambda(z)/16$ turns our algebraic power series $f(x)\in\mathbf{Z}\llbracket x\rrbracket$ into a boundary case (unit conformal size $\varphi^{\prime}(0)=1$) of André’s criterion. Another boundary case, but this time transcendental and incidentally demonstrating the sharpness of the qualitative André algebraicity criterion even in the a priori holonomic situation (see [And04, Appendix, A.5] for a discussion), is provided by the Gauss hypergeometric function
@@
which transforms this hypergeometric series into a weight one modular form for the congruence group $\Gamma_{0}(4)$.
@@ 1.1.3. A finite local monodromy leads to an overconvergence
It turns out, and this is the key to our method and already answers André’s question in [And04, Appendix, A.5], that a different choice of $\varphi(z)$ allows one to arithmetically distinguish between these two cases (algebraic and transcendental), and to have the algebraicity of $f(x)$ recognized by André’s Diophantine criterion by way of an “overconvergence.” The common feature of these two functions $f(x)$ and $F(x)$ — coming respectively out of modular forms of weights 0 and $1$ — is that they both vary holonomically in $x\in\mathbf{C}\smallsetminus\{0,1/16\}$: they satisfy linear ODEs with coefficients in $\mathbf{Q}[x]$ and no singularities apart from the three punctures $x=0,1/16,\infty$ of $Y(2)=\mathbf{H}/\Gamma(2)$. The difference feature is that their respective local monodromies around $x=0$ are finite for the case of $f(x)$ (a quotient of $\mathbf{Z}/N$, with the order $N$ equal to the LCM of cusp widths, or Wohlfahrt level [Woh64] of $f(x)$); and infinite for the case of $F(x)$ (isomorphic to $\mathbf{Z}$, corresponding more particularly to the fact that this particular hypergeometric function acquires a $\log{x}$ term after an analytic continuation around a small circle enclosing $x=1/16$). If now we perform the variable change $x\mapsto x^{N}$, redefaulting to $x:=\sqrt[N]{\lambda(q^{N})/16}$, that resolves the $N$-th root ambiguity in the formal Puiseux branches of $f(x)$ at $x=0$, and the resulting algebraic power series $f(x^{N})\in\mathbf{Z}\llbracket x^{N}\rrbracket\subset\mathbf{Z}\llbracket x\rrbracket$ has turned holonomic on $\mathbf{P}^{1}\smallsetminus\{16^{-1/N}\mu_{N},\infty\}$: singularities only at $16^{-1/N}\mu_{N}\cup\{\infty\}$ (but not at $x=0$: this key step of exploiting arithmetic algebraization is the same as in Ihara’s arithmetic connectedness theorem [Iha94, Theorem 1], which together with Bost’s extension [Bos99] to arithmetic Lefschetz theorems have in equal measure been inspirational for our whole approach to the unbounded denominators conjecture). Since $\lambda:D(0,1)\to\mathbf{C}\smallsetminus\{1\}$ has fiber $\lambda^{-1}(0)=\{0\}$, the function $\varphi(z):=\sqrt[N]{\lambda(z^{N})/16}:D(0,1)\to\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N}$ is still holomorphic on the unit disc $|z|<1$, and under this tautological choice, both functions $f(x^{N})$ and $F(x^{N})$ continue to be at the borderline of André’s algebraicity criterion: $|\varphi^{\prime}(0)|=1$.
@@
But if instead of the tautological simultaneous uniformization we take $\varphi:D(0,1)\to\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N}$ to be the universal covering map (pointed at $\varphi(0)=0$), then either by a direct computation with monodromy, or by Cauchy’s analyticity theorem on the solutions of linear ODEs with analytic coefficients and no singularities in a disc, we have both function germs $x:=\varphi(z)$ and $f(x):=f(\varphi(z))$ holomorphic, hence convergent, on the full unit disc $D(0,1)$. In contrast, now $F(\varphi(z))$ converges only up to the “first” nonzero fiber point in $\varphi^{-1}(0)\smallsetminus\{0\}$, giving a certain radius rather smaller than $1$. We must have the strict lower bound $|\varphi^{\prime}(0)|>1$, because the preceding unit-radius holomorphic map $\sqrt[N]{\lambda(z^{N})/16}:D(0,1)\to\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N}$ has to factorize properly through the universal covering map. Indeed in Theorem 5.3.8, using an explicit description by hypergeometric functions of the multivalued inverse of the universal covering map of $\mathbf{C}\smallsetminus\mu_{N}$ based on Poincaré’s ODE approach [Hem88] to the uniformization of Riemann surfaces, we find an exact formula for this uniformization radius in terms of the Euler Gamma function. [^1] André pointed out to us that this explicit formula has previously been obtained by Kraus and Roth, see [KR16, Remark 5.1]. Hence the algebraicity of $f(x)$ gets witnessed by André’s criterion; and the formal new result that we get already at this opening stage (see Theorem 7.2.1) is that any integral formal power series solution $f(x)\in\mathbf{Z}\llbracket x\rrbracket$ to a linear ODE $L(f)=0$ without singularities on $\mathbf{P}^{1}\smallsetminus\{0,1/16,\infty\}$ is in fact algebraic as soon as the linear differential operator $L$ has a finite local monodromy $\mathbf{Z}/N$ around the singular point $x=0$. More than this: the quantitative Theorem 2.0.2 proves that the totality of such $f(x)\in\mathbf{Z}\llbracket x\rrbracket$ at a given $N$ span a finite-dimensional $\mathbf{Q}(x)$-vector space, and gives an upper bound on its dimension as a function of the Wohlfahrt level parameter $N$. Now since a (noncongruence) counterexample $f(\tau)\in\mathbf{Z}\llbracket q\rrbracket$ to Theorem 1.0.1 would not exist on its own but entail a whole sequence $f(\tau/p)\in\mathbf{Z}\llbracket q^{1/p}\rrbracket$ of $\mathbf{Q}(x)$-linearly independent counterexamples at growing Wohlfahrt level $N\mapsto Np$, our idea is to measure up the supply of these putative (fictional) counterexamples alongside the congruence supply at a gradually increasing level until together they break the quantitative bound (2.0.4) supplied by our arithmetic holonomy Theorem 2.0.2.

### v2

@@ 1. Introduction
(1.0.2) $\frac{\lambda(\tau)}{16}=\left(\frac{\eta(\tau/2)\eta(2\tau)^{2}}{\eta(\tau)^{3}}\right)^{8}=q\prod_{n=1}^{\infty}\left(\frac{1+q^{2n}}{1+q^{2n-1}}\right)^{8}=q-8q^{2}+\cdots$
with $q=e^{\pi i\tau}$ and $\eta(\tau/2)=q^{1/24}\prod_{n=1}^{\infty}(1-q^{n})$. (Historic conventions force one to use $q$ for both $e^{\pi i\tau}$ and $e^{2\pi i\tau}$ — we use the first choice unless we expressly state otherwise.) On replacing the weight $k$ form $f$ by the weight zero form $f(\tau)(\lambda(\tau)/16\eta(\tau/2)^{2})^{k}$, we may (and do) assume that $k=0$. The function $f$ is then an algebraic function of $\lambda$, with branching only at the three punctures $\lambda=0,1,\infty$ of the modular curve $Y(2)\cong\mathbf{P}^{1}\smallsetminus\{0,1,\infty\}$. Thus another reading of our result states that the Belyĭ maps (étale coverings)
@@
The resulting Fourier coefficients are closely linked to the dimensions of the irreducible representations of the exceptional Lie group $E_{8}(\mathbf{C})$, and in particular they are integers. To be more precise: the modular function $j^{1/3}$ coincides with the graded dimension of the level one highest-weight representation of the affine Kac–Moody algebra $E_{8}^{(1)}$; see Gannon’s book [Gan06, page 196] for a broad view on this topic and its relation to mathematical physics. The unbounded denominators conjecture (Theorem 1.0.1) now implies that $j^{1/3}$ must be a modular function on a congruence subgroup, and indeed one may easily confirm that it is a Hauptmodul for the level $3$ subgroup which is the kernel of the composite $\mathrm{PSL}_{2}(\mathbf{Z})\rightarrow\mathrm{PSL}_{2}(\mathbf{F}_{3})=A_{4}\rightarrow\mathbf{Z}/3\mathbf{Z}$. One final example is the function
$h:=\frac{\lambda(\tau)(1-\lambda(\tau))}{16}=\left(\frac{\eta(\tau/2)\eta(2\tau)}{\eta(\tau)^{2}}\right)^{24}$
of level $\Gamma^{0}(2)\supset\Gamma(2)$; here the complete list of $n$ for which $h^{1/n}$ is either congruence modular or has bounded denominators are the divisors of $24$. One may even compute that the abelianization of the congruence completion $\widehat{\Gamma^{0}(2)}$ of $\Gamma^{0}(2)$ (where $\Gamma^{0}(2)$ is considered as a subgroup of $\mathrm{PSL}_{2}(\mathbf{Z})$) is $\mathbf{Z}/24\mathbf{Z}\oplus\mathbf{Z}/2\mathbf{Z}$. (The other $\mathbf{Z}/2\mathbf{Z}$ extension corresponds to $\sqrt{1-64h}=1-2\lambda$.)
In a similar vein pertaining to the examples from the representation theory of vertex operator algebras, we prove in our closing section 7 the natural generalization of Theorem 1.0.1 to components of vector-valued modular forms for $\mathrm{SL}_{2}(\mathbf{Z})$, in particular resolving — in a sharper form, in fact — Mason’s unbounded denominators conjecture [Mas12, KM08] on generalized modular forms.
@@ 1.1.1. The Diophantine principle
We shall refer to the first inequality as a Liouville lower bound, following its use by Liouville in his proof of the lower bound $|\alpha-p/q|\gg 1/q^{n}$ for algebraic numbers $\alpha\neq p/q$ of degree $n\geq 1$. We shall refer to the second inequality as a Cauchy upper bound, following the example above where it comes from an application of the Cauchy integral formula. The first non-trivial generalization of Lemma 1.1.1 was Émile Borel’s theorem [Bor94]. Dwork famously used a $p$-adic generalization of Borel’s theorem in his $p$-adic analytic proof of the rationality of the zeta function of an algebraic variety over a finite field (see Dwork’s account in the book [DGS94, Chapter 2]). The simplest non-trivial statement of Borel’s theorem is that an integral formal power series $f(x)\in\mathbf{Z}\llbracket x\rrbracket$ must already be a rational function as soon as it has a meromorphic representation as a quotient of two convergent complex-coefficients power series on some disc $D(0,R)$ of a radius $R>1$. The subject of arithmetic algebraization blossomed at the hands of many authors, including most prominently Carlson, Pólya, Robinson, Salem, Cantor, D. &\& G. Chudnovsky, Bertrandias, Zaharjuta, André, Bost, Chambert-Loir [CL02, BCL09, Ami75], [And04, § I.5], [And89, § VIII]. A simple milestone that we further develop in our § 2 is André’s algebraicity criterion [And04, Théorème 5.4.3], stating in a particular case that an integral formal power series $f(x)\in\mathbf{Z}\llbracket x\rrbracket$ is algebraic as soon as the two formal functions $x$ and $f$ admit a simultaneous analytic uniformization — that means an analytic map $\varphi:(D(0,1),0)\to(\mathbf{C},0)$ such that the composed formal function germ $f(\varphi(z))\in\mathbf{C}\llbracket z\rrbracket$ is also analytic on the full disc $D(0,1)$, and such that $\varphi$ is sufficiently large in terms of conformal size, namely: $|\varphi^{\prime}(0)|>1$. For example, for any integer $m$, the algebraic power series $f=(1-m^{2}x)^{1/m}\in\mathbf{Z}\llbracket x\rrbracket$ admits the simultaneous analytic uniformization $x=(1-e^{Mz})m^{-2}$ and $f=e^{Mz/m}$, where the conformal size $|\varphi^{\prime}(0)|=M/m^{2}$ can clearly be made arbitrarily large by making a suitable choice of $M$.
@@
In our setting, the Liouville lower bound ultimately comes down to the integrality (“bounded denominators”) hypothesis on the Fourier coefficients of $f(\tau)$, while the Cauchy upper bound comes down to studying the mean growth behavior $m(r,\varphi):=\int_{|z|=r}\log^{+}{|\varphi|}\,\mu_{\mathrm{Haar}}$ of the largest (universal covering) analytic map $\varphi:D(0,1)\to\mathbf{C}\smallsetminus\mu_{N}$ avoiding the $N$-th roots of unity. These are clearly distinguished in our abstract arithmetic algebraization work of § 2 as the steps (2.2.2) and (2.2.3), respectively. (We also refer to (2.4.8) and (2.4.2), resp. (2.5.23) and (2.5.22), in our alternative treatments.) Our Theorem 2.0.1 is effectively a quantitative refinement of André’s algebraicity criterion to take into account the degree of algebraicity over $\mathbf{Q}(x)$, and still more precisely a certain holonomy rank over $\mathbf{Q}(x)$. Foreshadowing a key technical point (to be discussed in more detail later in the introduction), our Cauchy upper bound is given in terms of a mean (integrated) growth term rather than a supremum term, and this improvement is essential to our approach.
@@ 1.1.2. Modularity and simultaneous uniformizations of $f$ and $\lambda$
of (1.0.2) has integer coefficients, expressing the identity $\mathbf{Z}\llbracket q\rrbracket=\mathbf{Z}\llbracket x\rrbracket$ of formal power series rings, and that formal substitution turns our integral Fourier coefficients hypothesis into an algebraic power series with integer coefficients: henceforth in this introductory sketch we switch to writing, by a mild and harmless notational abuse, simply $f(x)\in\mathbf{Z}\llbracket x\rrbracket$ in place of $f(\tau)$ and $\lambda(q)$ in place of (1.0.2). In the general case of arbitrary cusp width, which we need anyhow for the inner workings of our proof even if one is ultimately interested in the $\mathbf{Z}\llbracket q\rrbracket$ case, we will only have $f\in\mathbf{Z}[1/N]\llbracket x\rrbracket$ when we write out $f(\tau)$ as a power series in $x:=\sqrt[N]{\lambda/16}$ to accommodate the Puiseux series — but there is still a hidden integrality property which we can exploit. That leads to some mild technical nuance with the power series (2.0.2) — think of $t=q^{1/N}$, $x(t)=\sqrt[N]{\lambda(t^{N})/16}$ and $p(x)=x^{N}$ — in our refinement (2.0.3) of André’s theorem.
The complex analysis enters by way of a linear ODE in the following way. To start with, we have, just by fiat, the simultaneous analytic uniformization of the two functions $x:=\lambda/16$ and $f$ by the complex unit $q$-disc $|q|<1$. In this way, the tautological choice $\varphi(z):=\lambda(z)/16$ turns our algebraic power series $f(x)\in\mathbf{Z}\llbracket x\rrbracket$ into a boundary case (unit conformal size $\varphi^{\prime}(0)=1$) of André’s criterion. Another boundary case, but this time transcendental and incidentally demonstrating the sharpness of the qualitative André algebraicity criterion even in the a priori holonomic situation (see [And04, Appendix, A.5] for a discussion), is provided by the Gauss hypergeometric function
@@
which transforms this hypergeometric series into a weight one modular form for the congruence group $\Gamma(2)$. The existence of such transcendental $\mathbf{Z}\llbracket x\rrbracket$ holonomic functions on $\mathbf{C}\smallsetminus\{0,1/16\}$ recovers—by André’s algebraicity criterion—a classical “$1/16$ theorem” of Carathéodory [Car54, (412.8) on page 198]. (See also Goluzin [Gol69, § III.1, Theorem 1].)
@@ 1.1.3. A finite local monodromy leads to an overconvergence
It turns out, and this is the key to our method and already answers André’s question in [And04, Appendix, A.5], that a different choice of $\varphi(z)$ allows one to arithmetically distinguish between these two cases (algebraic and transcendental), and to have the algebraicity of $f(x)$ recognized by André’s Diophantine criterion by way of an “overconvergence.” The common feature of these two functions $f(x)$ and $F(x)$ — coming respectively out of modular forms of weights 0 and $1$ — is that they both vary holonomically in $x\in\mathbf{C}\smallsetminus\{0,1/16\}$: they satisfy linear ODEs with coefficients in $\mathbf{Q}[x]$ and no singularities [^1] With nontrivial local monodromy. The precise definition is in 2.0.4. apart from the three punctures $x=0,1/16,\infty$ of $Y(2)=\mathbf{H}/\Gamma(2)$. The difference feature is that their respective local monodromies around $x=0$ are finite for the case of $f(x)$ (a quotient of $\mathbf{Z}/N$, with the order $N$ equal to the LCM of cusp widths, or Wohlfahrt level [Woh64] of $f(x)$); and infinite for the case of $F(x)$ (isomorphic to $\mathbf{Z}$, corresponding more particularly to the fact that this particular hypergeometric function acquires a $\log{x}$ term after an analytic continuation around a small circle enclosing $x=1/16$). If now we perform the variable change $x\mapsto x^{N}$, redefaulting to $x:=\sqrt[N]{\lambda(q^{N})/16}$, that resolves the $N$-th root ambiguity in the formal Puiseux branches of $f(x)$ at $x=0$, and the resulting algebraic power series $f(x^{N})\in\mathbf{Z}\llbracket x^{N}\rrbracket\subset\mathbf{Z}\llbracket x\rrbracket$ has turned holonomic on $\mathbf{P}^{1}\smallsetminus\{16^{-1/N}\mu_{N},\infty\}$: singularities only at $16^{-1/N}\mu_{N}\cup\{\infty\}$ (but not at $x=0$: this key step of exploiting arithmetic algebraization is the same as in Ihara’s arithmetic connectedness theorem [Iha94, Theorem 1], which together with Bost’s extension [Bos99] to arithmetic Lefschetz theorems have in equal measure been inspirational for our whole approach to the unbounded denominators conjecture). Since $\lambda:D(0,1)\to\mathbf{C}\smallsetminus\{1\}$ has fiber $\lambda^{-1}(0)=\{0\}$, the function $\varphi(z):=\sqrt[N]{\lambda(z^{N})/16}:D(0,1)\to\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N}$ is still holomorphic on the unit disc $|z|<1$, and under this tautological choice, both functions $f(x^{N})$ and $F(x^{N})$ continue to be at the borderline of André’s algebraicity criterion: $|\varphi^{\prime}(0)|=1$.
@@
But if instead of the tautological simultaneous uniformization we take $\varphi:D(0,1)\to\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N}$ to be the universal covering map (pointed at $\varphi(0)=0$), then either by a direct computation with monodromy, or by Cauchy’s analyticity theorem on the solutions of linear ODEs with analytic coefficients and no singularities in a disc, we have both function germs $x:=\varphi(z)$ and $f(x):=f(\varphi(z))$ holomorphic, hence convergent, on the full unit disc $D(0,1)$. In contrast, now $F(\varphi(z))$ converges only up to the “first” nonzero fiber point in $\varphi^{-1}(0)\smallsetminus\{0\}$, giving a certain radius rather smaller than $1$. We must have the strict lower bound $|\varphi^{\prime}(0)|>1$, because the preceding unit-radius holomorphic map $\sqrt[N]{\lambda(z^{N})/16}:D(0,1)\to\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N}$ has to factorize properly through the universal covering map. Indeed in Theorem 5.1.4, using an explicit description by hypergeometric functions of the multivalued inverse of the universal covering map of $\mathbf{C}\smallsetminus\mu_{N}$ based on Poincaré’s ODE approach [Hem88] to the uniformization of Riemann surfaces, we find an exact formula for this uniformization radius in terms of the Euler Gamma function. [^2] André pointed out to us that this explicit formula has previously been obtained by Kraus and Roth, see [KR16, Remark 5.1]. See also [Gol69, §III.1]. Hence the algebraicity of $f(x)$ gets witnessed by André’s criterion; and the formal new result that we get already at this opening stage (see Theorem 7.2.1) is that any integral formal power series solution $f(x)\in\mathbf{Z}\llbracket x\rrbracket$ to a linear ODE $L(f)=0$ without singularities on $\mathbf{P}^{1}\smallsetminus\{0,1/16,\infty\}$ is in fact algebraic as soon as the linear differential operator $L$ has a finite local monodromy $\mathbf{Z}/N$ around the singular point $x=0$. More than this: the quantitative Corollary 2.0.5 proves that the totality of such $f(x)\in\mathbf{Z}\llbracket x\rrbracket$ at a given $N$ span a finite-dimensional $\mathbf{Q}(x)$-vector space, and gives an upper bound on its dimension as a function of the Wohlfahrt level parameter $N$. Now since a (noncongruence) counterexample $f(\tau)\in\mathbf{Z}\llbracket q\rrbracket$ to Theorem 1.0.1 would not exist on its own but spawn a whole sequence $f(\tau/p)\in\mathbf{Z}\llbracket q^{1/p}\rrbracket$ of $\mathbf{Q}(x)$-linearly independent counterexamples at growing Wohlfahrt level $N\mapsto Np$, our idea is to measure up the supply of these putative (fictional) counterexamples alongside the congruence supply at a gradually increasing level until together they break the quantitative bound (2.0.3) supplied by our arithmetic holonomy Theorem 2.0.1.